[Brahms-l] Shift report 20050128 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jan 28 2005 - 15:38:18 EST
1.  The machine has been down for all but the fist 1.5 hours of this shift.  

2.  I took the time to take some silicon pulser calibration runs.

3. Hiro and Bob have been working on replacing a camac crate for the silicon.  

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jan 28 15:38:37 2005

This archive was generated by hypermail 2.1.8 : Fri Jan 28 2005 - 15:39:02 EST